OpenStudy (anonymous):

Solve for x: 3(x + 1) = −2(x − 1) − 4

OpenStudy (luigi0210):

Distribute first

OpenStudy (anonymous):

3(x+1)=-2(x-1)-4 =>3x+3=-2x+2-4 =>3x+2x=-2-3 =>5x=-5 x=-1

OpenStudy (luigi0210):

You can't just give answers like that @SidK you have to lead them to it..

OpenStudy (anonymous):

fist of all i am leading them!and everyone does the same!

OpenStudy (luigi0210):

Yes, but most others give an explanation. Not just the answer

OpenStudy (anonymous):

i am not giving the answer only i tried to solve it !and solved !i know it would be easy for him/her to understand!

OpenStudy (luigi0210):

Let's hope so.

OpenStudy (anonymous):

yeah!